OpenStudy (anonymous):

Find a polynomial of degree with real coefficients and the following zeros: 4,2-i

OpenStudy (amoodarya):

when you have real coefficients and one of root is 4-i another root must be 4+i so can you go on ?

OpenStudy (amoodarya):

like this suppose one of root is 3+2i with real coefficient f(x)=a(x-(3+2i))(x-(3-2i) ) = a(x^2 +(9-(2i)^2- 2x(3+2i+3-2i)) =a(x^2+(9+4)+2x(6))